A nuclear reactor continuously generates 150 MW of power through the fissioning of uranium. Suppose that each fission releases 190 MeV. If one mole of uranium (6.023 ×10^23 nuclei) has a mass of 0.235 kg, what mass of uranium has undergone fission in a 4.0 day period?
1
Expert's answer
2013-01-09T08:18:33-0500
Let us count number of nucleus(atoms)which had decayed:
4 days * 150 MW/(190 MeV)& = 4*86,400s * 150*10^6 J/s /(3*10^(-11) J) = 1.728 * 10^(24)

Let us find their mass:
0.235 kg * 1.728*10^(24)/(6.023*10^(23)) = 674.21 g = 0.67421 kg